# Retrieval-Augmented Generation in Biomedicine: A Survey of Technologies, Datasets, and Clinical Applications

### v2

@@ Introduction
MedMCQA 209 2022 Multiple-choice Self-BioRAG 106, JMLR 3, Ascle 30, Liévin et al. 206, EyeGPT 153, PodGPT 34
ChatDoctor 210 2023 Dialogue EyeGPT 153
MedAlpaca 173 2023 Dialogue EyeGPT 153
GeneTuring 211 2023 Genomics QA GeneGPT 199
GIT 212 2023 Information Extraction BiomedRAG 4
MIRAGE 61 2024 Multiple-choice & QA MEDRAG 61
MedExpQA 51 2024 Multilingual Medical QA MedExpQA 51, PodGPT 34
PubMedRS-200 14 2024 Medical QA Clinifo.ai 14
MedicineQA 182 2024 Multi-round dialogue RagPULSE 182
CELLS 27 2024 Lay language generation Guo 27
Applications
@@
The application of RAG in the biomedical domain represents a significant advancement in healthcare informatics, combining the strengths of LLMs with domain-specific knowledge retrieval. This integration helps mitigate challenges inherent to healthcare applications—factual accuracy, domain specificity, knowledge recency, and explainability—that conventional generative AI approaches struggle to overcome 213. In this section, we analyze key healthcare domains where biomedical RAG systems have demonstrated substantial utility and are organized according to their functional contributions to clinical practice, precision medicine, and healthcare knowledge dissemination.
@@
Medical question-answering systems provide clinicians and patients with access to accurate healthcare information at the point of need 100. This application is essential because it enables rapid retrieval of evidence-based information, supporting clinical decision-making while reducing the cognitive load on healthcare providers and improving information access for patients with varying levels of health literacy.
Recent advancements have produced several sophisticated medical QA systems with distinctive approaches to knowledge retrieval. Clinfo.ai 14 provides evidence-based healthcare QA by retrieving information from trusted sources like PubMed using the Entrez API, ensuring responses reflect current medical evidence. RALL 27 employs Dense Passage Retriever to generate lay language explanations of medical concepts, addressing the critical need for accessible medical information. Self-BioRAG 106 implements the MedCPT retriever with self-reflection techniques to improve accuracy for complex clinical questions with multiple components, while WeiseEule 101 integrates multiple retrievers (MedCPT, BioBERT, BioGPT) with Pinecone DB to handle diverse medical information needs across specialties. These medical QA systems collectively demonstrate RAG’s potential for healthcare information access, providing contextually relevant, evidence-based responses to complex medical queries for both clinical decision support and patient education.
@@
Diagnostic and treatment decision support systems assist clinicians in navigating complex clinical pathways through evidence retrieval and synthesis 214. They can help healthcare providers identify relevant evidence, recognize patterns, and identify appropriate interventions, ultimately improving diagnostic accuracy and treatment selection while reducing cognitive biases and diagnostic errors.
Several innovative implementations showcase RAG’s potential in clinical decision support across medical specialties. Expert-Guided LLMs 15 support specialized decision-making in precision oncology by retrieving relevant genomic information and matching it with treatment guidelines and clinical trial data. DR.KNOWS 17 facilitates diagnostic prediction by connecting EHRs with medical knowledge repositories through knowledge graph-enhanced retrieval that models complex relationships between symptoms, diseases, and treatments. CliniqIR 52 combines MedCPT and BM25 retrieval to match patient symptoms with potential diagnoses, while SurgeryLLM 28 delivers surgical decision support by retrieving relevant procedures, techniques, and complication data. RECTIFIER 91 employs OpenAI Text-Embedding with Faiss for clinical trial screening, helping match heart failure patients with appropriate clinical trials by analyzing eligibility criteria against patient data. By connecting patient-specific data with evidence-based medical knowledge, these RAG systems significantly enhance clinical decision-making while reducing the cognitive burden on healthcare providers, potentially improving both efficiency and quality of care.
@@
Rare disease identification and management represents a particularly challenging healthcare domain due to limited literature, dispersed expertise, and complex symptom patterns 215. RAG systems are especially valuable in this context because they can aggregate and synthesize knowledge from diverse sources, helping clinicians recognize unusual patterns and connect them with appropriate diagnostic and management strategies for conditions they may encounter only rarely in clinical practice.
Recent research has produced several specialized rare disease applications leveraging different knowledge retrieval approaches. Zelin et al. 16 developed a RAG system specifically focused on rare disease diagnosis, connecting uncommon symptom patterns with rare disease knowledge drawn from specialized databases. RAG-HPO 72 utilizes FastEmbed for automated deep phenotyping of rare genetic disorders, linking phenotypic observations with potential genetic causes through human phenotype ontology mapping. Yang et al. 216 created RDguru, an intelligent conversational agent that employs a multi-source fusion diagnostic model combining GPT-4, PheLR, and phenotype matching strategies to significantly improve diagnostic recall for rare diseases. Some authors 217 have also evaluated the capabilities of LLMs and RAG for phenotype-driven gene prioritization in rare genetic disease diagnosis. These applications demonstrate how RAG systems can effectively bridge knowledge gaps in rare disease management, connecting dispersed information to support timely diagnosis and appropriate treatment strategies for conditions where clinical expertise is often limited.
@@
Automatic and semi-automatic report generation can help healthcare providers by significantly reducing the time they spend on documentation, allowing them to dedicate more time to patient care 218. A critical area of report generation is radiology, which involves image analysis, document consultation, and data evaluation 219. This process faces multiple challenges including ensuring factual accuracy, maintaining clinical relevance, minimizing hallucinations, and providing sufficient interpretability for clinical adoption 220.
RAG frameworks have been proposed to address these challenges by grounding language generation in relevant retrieved content, resulting in more accurate and trustworthy outputs 7. Recent works have demonstrated the effectiveness of RAG in this domain. For example, multimodally-aligned embeddings with general-domain LLMs such as GPT-4 221, ensuring relevant content retrieval and mitigating hallucinations while allowing customization based on clinical intent. To enhance interpretability, some authors 222 proposed a multi-agent RAG framework utilizing concept bottleneck models that guide report generation with clinically meaningful concept vectors, bridging the gap between performance and explainability. RAG pipelines 223 have also been used to improve image-text matching techniques, incorporating semantic segmentation and triple contrastive loss, leading to more accurate and fluent reports with reduced errors. Differently, LaB-RAG 224 employs label-boosted retrieval by converting medical images into descriptive text labels, which are then used for text-based retrieval and report generation, bypassing the need for direct image processing by the LLM. Lastly, some authors 225 have proposed fact-aware multimodal retrieval strategies to address factual inconsistencies in generated reports by leveraging RadGraph-based factual mining and multimodal retrievers, ensuring high-quality reference retrieval and improved factual completeness.
@@
Genomic medicine involves analyzing genetic data to inform clinical decision-making. The rapid knowledge evolution and inherent complexity in variant interpretation within genomic medicine impose significant cognitive burdens on clinicians, particularly given the exponential expansion of genomic information and the challenges in ascertaining variant pathogenicity 226.
Several innovative RAG applications address these challenges in genomic medicine through specialized knowledge retrieval. FAVOR-GPT 87 utilizes OpenAI Text-Embedding with Weaviate DB to facilitate genome variant annotations, connecting genetic variants with clinical significance and supporting evidence from multiple databases. Lu et al. 18 developed a genomics analysis system using Azure AI Search with keyword matching for more accurate genetic variant interpretation within clinical contexts. RAG-HPO 72 connects clinical observations with genetic variants in rare disorders through sophisticated knowledge mapping, while JGCLLM 81 implements GLuCoSE-base-ja vectors to provide genetic counseling support that translates complex genomic findings into clinically meaningful information. By connecting genomic findings with clinical knowledge through sophisticated retrieval mechanisms, these RAG systems significantly improve the clinical application of genetic information, such as disease inference 18 and phenotypic and genotypic analysis 72, while supporting more personalized treatment decisions in the emerging era of genomic medicine.
@@
Personalized health management, a tailored healthcare approach, involves generating individualized health information and recommendations based on patient-specific characteristics, preferences, and health status 227. This application mitigates a important healthcare challenge: the need for more effective and patient-centered care that optimizes treatment adherence, self-management, and health outcomes, in contrast to traditional standardized approaches that often fail to accommodate specific patient needs, literacy levels, and medical conditions.
Recent research has produced several personalized health management systems employing different knowledge retrieval strategies. HEALIE 29 leverages knowledge graphs to generate personalized medical content incorporating both clinical knowledge and socio-economic factors relevant to individual patients. RISE 85 implements OpenAI Text-Embedding with FAISS to provide personalized responses to diabetes-related inquiries, helping patients navigate complex self-management requirements. A Dual RAG System 50 combines dense retrieval and BM25 with language-specific tokenizers to support diabetes management across multiple languages, expanding access to personalized health information for diverse populations. RAP 92 utilizes Amazon Titan Text Embeddings to provide personalized nutrition-related question answering, supporting dietary management for various health conditions. These personalized health management systems demonstrate RAG’s ability to bridge the gap between complex medical information and individual patient needs through contextually relevant knowledge retrieval, ultimately supporting better self-management and improved health outcomes through personalized guidance.
@@
Medical education and training systems support the development of healthcare professionals through evidence-based learning resources and simulated clinical experiences 228. These systems help address workforce needs by enabling more efficient knowledge acquisition, supporting the development of clinical reasoning skills, and allowing for personalized learning experiences that adapt to individual learner needs and knowledge gaps.
ChatENT 89 leverages Retrieval-Augmented Language Modeling (RALM) to support medical education in Otolaryngology-Head and Neck Surgery (OHNS). EyeGPT 153 transforms a general LLM into an ophthalmology-specialized assistant through prompt engineering, fine-tuning with eye-specific terminology, and RAG from ophthalmology textbooks, supporting both medical education and patient inquiries. Sevgi et al. 186 developed targeted applications in ophthalmology , including "EyeTeacher" for generating educational questions and "EyeAssistant" for clinical support, both enhanced through the integration of clinical guidelines and peer-reviewed documents for retrieval. In nursing education, Zhao et al 229 leveraged RAG and semantic search technology to answer questions from the Chinese Nursing Licensing Exam, showing the potential for specialized knowledge retrieval in professional certification. By connecting learning objectives with relevant medical knowledge through sophisticated retrieval mechanisms, these RAG-based education systems serve as valuable educational aid tools for healthcare practice skill training.
@@
Clinical research application tools are designed to support clinical studies by enhancing efficiency, leveraging knowledge retrieval and synthesis to accelerate the development of medical knowledge and therapeutic innovations. 230.

### v3

@@ Introduction
MedMCQA pal2022medmcqa 2022 Multiple-choice Self-BioRAG jeong2024improving, JMLR Wang2024JMLR, Ascle Ascle, Liévin et al. Lievin2024, EyeGPT chen2024eyegpt, PodGPT jia2024podgpt
ChatDoctor li2023chatdoctor 2023 Dialogue EyeGPT chen2024eyegpt
MedAlpaca han2023medalpaca 2023 Dialogue EyeGPT chen2024eyegpt
GeneTuring hou2023geneturing 2023 Genomics QA GeneGPT jin2024genegpt
GIT li2023petailor 2023 Information Extraction BiomedRAG li2025biomedrag
MIRAGE xiong-etal-2024-benchmarking 2024 Multiple-choice & QA MEDRAG xiong-etal-2024-benchmarking
MedExpQA ALONSO2024102938 2024 Multilingual Medical QA MedExpQA ALONSO2024102938, PodGPT jia2024podgpt
PubMedRS-200 Lozano2024ClinfoAI 2024 Medical QA Clinifo.ai Lozano2024ClinfoAI
MedicineQA huang2024toolcalling 2024 Multi-round dialogue RagPULSE huang2024toolcalling
CELLS Guo2024 2024 Lay language generation Guo Guo2024
Evaluation Metrics
The biomedical RAG systems reviewed in this survey typically follow a three-stage pipeline, consisting of (1) retrieval, (2) reranking, and (3) generation. Consequently, the evaluation methodologies used across the surveyed studies can also be naturally organized according to these stages. Because each stage involves a distinct functional component of the RAG architecture, the choice of evaluation metrics naturally varies across stages. Moreover, as shown in Table LABEL:tab:biomedical-comprehensive, biomedical RAG systems are applied to a wide range of downstream tasks—such as NER, medical QA, dialogue, and clinical summarization—and the selection of evaluation metrics further depends on the specific characteristics of these tasks. In what follows, we review the evaluation metrics used in biomedical RAG research by following this pipeline-based structure.
Metrics for Retrieval
The retrieval stage evaluates how effectively a system surfaces and orders clinically relevant documents from large-scale biomedical corpora, with emphasis on ranking quality, relevance precision, and coverage of essential evidence. More broadly, evaluation in this stage follows an information-retrieval paradigm, where the goal is to quantify how well the retrieved set and its ordering align with the ground-truth relevance distribution. We review the core retrieval metrics commonly adopted in biomedical RAG systems—namely, Precision@k, Recall@k, Mean Average Precision (MAP@k), and normalized Discounted Cumulative Gain (nDCG@k)—and provide their formal definitions.
  1. (1)
    Precision@k
    $\text{Precision@k}=\frac{\sum_{i=1}^{k}\mathbf{1}[\text{doc}_{i}\text{ is relevant}]}{k}$ (7)
  2. (2)
    Recall@k
    $\text{Recall@k}=\frac{\sum_{i=1}^{k}\mathbf{1}[\text{doc}_{i}\text{ is relevant}]}{\lvert\text{Relevant Documents}\rvert}$ (8)
  3. (3)
    Mean Average Precision (MAP@k)
    $\text{AP@k}=\frac{1}{R}\sum_{i=1}^{k}\text{Precision@i}\cdot\mathbf{1}[\text{doc}_{i}\text{ is relevant}]$ (9)
    $\text{MAP@k}=\frac{1}{N}\sum_{j=1}^{N}\text{AP@k}(j)$ (10)
  4. (4)
    Normalized Discounted Cumulative Gain (nDCG@k)
    $\text{DCG@k}=\sum_{i=1}^{k}\frac{2^{rel_{i}}-1}{\log_{2}(i+1)}$ (11)
    $\text{nDCG@k}=\frac{\text{DCG@k}}{\text{IDCG@k}}$ (12)
Metrics for Reranking
The reranking stage refines the ordering of retrieved documents using more expressive scoring mechanisms (e.g., cross-encoder relevance models). Because its objective remains aligned with retrieval—prioritizing the most clinically relevant evidence—biomedical RAG systems typically evaluate reranking with the same IR metrics: Precision@k, Recall@k, MAP@k, and nDCG@k.Notably, in RAG applications, reranking quality is particularly significant because even small improvements in the top-ranked documents can substantially enhance the factual grounding and reliability of the final generated output.
Metrics for Generation
The generation stage is evaluated primarily through task-specific criteria, as the correctness and quality of the final output depend heavily on the nature of the downstream biomedical task. Because different applications—such as medical QA, clinical summarization, diagnostic classification, and dialogue—impose distinct requirements, the choice of evaluation metrics must be aligned with the characteristics of each task. In what follows, we group representative evaluation metrics into four categories: (1) generation metrics; (2) classification metrics; (3) named entity recognition metrics for biomedical IE tasks; and (4) dialogue metrics for assessing multi-turn clinical conversations.
Generation Metrics assess the quality of free-form textual outputs by measuring aspects such as content correctness, semantic alignment, and linguistic coherence. They are applicable to downstream biomedical generation tasks that require producing narrative responses, including medical QA, clinical report generation, and clinical summarization.
  1. (1)
    ROUGE-N
    $\text{ROUGE-N}=\frac{\text{Overlap of N-grams}}{\text{Total N-grams in reference}}$ (13)
    ROUGE-N measures the n-gram overlap between generated text and the reference text, commonly used for summarization and long-form generation.
  2. (2)
    ROUGE-L
    $\text{ROUGE-L}=\frac{\text{LCS}(X,Y)}{\text{Length of reference}}$ (14)
    ROUGE-L is based on the Longest Common Subsequence (LCS), capturing sentence-level structural similarity beyond exact n-gram matching.
  3. (3)
    BLEU
    $\text{BLEU}=\text{BP}\cdot\exp\left(\sum_{n=1}^{4}w_{n}\log p_{n}\right)$ (15)
    BLEU evaluates the precision of n-grams and is widely adopted for machine translation and other structured text generation tasks.
  4. (4)
    BERTScore
    BERTScorebert-score evaluates semantic similarity between a generated text $X$ and a reference text $Y$ by aligning their token embeddings from a pretrained language model. Let $e_{x}$ and $e_{y}$ denote contextual embeddings of tokens $x\in X$ and $y\in Y$, respectively. The precision $P$ and recall $R$ are defined as
    $P=\frac{1}{|X|}\sum_{x\in X}\max_{y\in Y}\cos\bigl(e_{x},e_{y}\bigr),\quad R=\frac{1}{|Y|}\sum_{y\in Y}\max_{x\in X}\cos\bigl(e_{y},e_{x}\bigr),$ (16)
    and the final BERTScore is computed as the $F_{1}$ score:
    $\text{BERTScore}_{F_{1}}=\frac{2\cdot P\cdot R}{P+R}.$ (17)
    Compared to lexical overlap metrics, BERTScore offers a more contextualized evaluation by operating in the embedding space.
  5. (5)
    BLEURT
    BLEURTsellam-etal-2020-bleurt is a learned evaluation metric based on pretrained language models fine-tuned on human-rated text pairs. Given a generated text $X$ and reference text $Y$, BLEURT encodes both sequences with a transformer encoder and predicts a scalar quality score:
    $\text{BLEURT}(X,Y)=f_{\theta}\bigl(\text{Enc}(X),\,\text{Enc}(Y)\bigr),$ (18)
    where $\text{Enc}(\cdot)$ denotes contextual embeddings and $f_{\theta}$ is a regression head trained to approximate human judgments. BLEURT captures semantic adequacy and fluency beyond surface-level overlap.
  6. (6)
    RaTEScore
    RaTEScorezhao2024ratescore is an entity-based metric for radiology report generation. It extracts medical entities and their relations from both generated and reference reports, matches them with synonym awareness, weights them by clinical importance, and computes an F1-style score. This makes the metric focus on factual, clinically meaningful consistency rather than surface wording, achieving higher alignment with expert judgment.
Classification metrics evaluate the correctness of discrete predictions made by a model, such as assigning labels or choosing among predefined options. These metrics are well-suited for tasks like diagnosis classification and multiple-choice question answering.
  1. (1)
    Accuracy (ACC) measures the proportion of correctly predicted instances:
    $\text{ACC}=\frac{\text{\# correct predictions}}{\text{\# total instances}}$ (19)
  2. (2)
    Precision, Recall, and F1 for a single class are defined as:
    $\text{Precision}=\frac{TP}{TP+FP},\quad\text{Recall}=\frac{TP}{TP+FN},\quad\text{F1}=\frac{2\cdot\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$ (20)
  3. (3)
    Macro-F1 is commonly used in multi-class and multi-label classification settings. It computes the F1 score independently for each class and then averages them, giving equal weight to both frequent and rare classes:
    $\text{Macro-F1}=\frac{1}{C}\sum_{c=1}^{C}\frac{2\cdot\text{Precision}_{c}\cdot\text{Recall}_{c}}{\text{Precision}_{c}+\text{Recall}_{c}}$ (21)
  4. (4)
    Micro-F1 is another metric for multi-class and multi-label classification tasks. It aggregates true positives, false positives, and false negatives across all classes before computing a global F1 score, thus favoring frequent classes:
    $\text{Micro-F1}=\frac{2\cdot\sum_{c=1}^{C}TP_{c}}{2\cdot\sum_{c=1}^{C}TP_{c}+\sum_{c=1}^{C}FP_{c}+\sum_{c=1}^{C}FN_{c}}$ (22)
AUROC and AUPRC are probability-based evaluation metrics commonly used to assess the performance of binary classifiers across different decision thresholds. We notice that they are rarely reported in the surveyed works, as LLMs typically generate free-text outputs or are prompted to select a single label from a predefined set rather than produce explicit class probabilities.
Named Entity Recognition (NER) metrics measures how accurately a system identifies and labels biomedical entities in text, and are widely used for assessing biomedical information extraction tasks.
  1. (1)
    Token-level F1 evaluates NER performance at the token granularity by computing the F1 score over token-wise precision and recall:
    $\text{Token-F1}=\frac{2\cdot P\cdot R}{P+R}$ (23)
  2. (2)
    Entity-level F1 assesses how accurately complete entity spans are identified, considering correct entities, boundary errors, and missed entities:
    $\text{Entity-F1}=\frac{2\cdot\lvert\text{Correct Entities}\rvert}{2\cdot\lvert\text{Correct Entities}\rvert+\lvert\text{Wrong Boundaries}\rvert+\lvert\text{Missed Entities}\rvert}$ (24)
Medical Dialogue evaluation is inherently complex, as it must account for both the linguistic quality of generated responses and the effectiveness of completing dialogue-level clinical tasks. In practice, evaluation combines automatic metrics with human or LLM-based judgments, since purely automated measures are often insufficient for safety-sensitive biomedical scenarios.
  1. (1)
    Generation quality (automatic): Metrics such as ROUGE, BLEU, BERTScore, and DialogRPTgao-etal-2020-dialogue are used to automatically assess lexical overlap, semantic similarity, and response coherence. These metrics provide scalable proxies for linguistic quality but do not capture clinical correctness.
  2. (2)
    Task success (primarily automatic): Indicators such as task completion rateli2021how and slot accuracyaksu2024granular measure whether the system achieves predefined clinical or informational goals. These metrics can be computed automatically when gold annotations are available, although human assessment is sometimes required for semantically defined success criteria.
  3. (3)
    Safety and factuality (human/LLM-based): Evaluation of hallucination, factual consistency, and clinical appropriateness typically relies on expert reviewers or medically aligned LLM judges Croxford2025, metric-survey, as automated approaches have limited ability to detect subtle or domain-specific medical errors. Automatic hallucination or consistency scores may be used as auxiliary signals, but do not replace human or expert-guided evaluation.
Together, these complementary evaluation strategies constitute a diverse assessment framework, and surveyed studies typically adopt one or more of these metrics to evaluate the quality of RAG-based medical dialogue systems.
Applications in biomedicine
@@
The application of RAG in the biomedical domain represents a significant advancement in healthcare informatics, combining the strengths of LLMs with domain-specific knowledge retrieval. This integration helps mitigate challenges inherent to healthcare applications—factual accuracy, domain specificity, knowledge recency, and explainability—that conventional generative AI approaches struggle to overcome liu2025improving. In this section, we analyze 5 key domains in which we identified that biomedical RAG systems have demonstrated substantial utility: clinical decision support, medical documentation, precision medicine, education, and clinical research.
@@
Medical question-answering systems provide clinicians and patients with access to accurate healthcare information at the point of need Soman2024. This application is particularly valuable because it enables rapid retrieval of evidence-based information, supporting clinical decision-making while reducing the cognitive load on healthcare providers and improving information access for patients with varying levels of health literacy.
Recent advancements have produced several novel medical QA systems with distinctive approaches to knowledge retrieval. Clinfo.ai Lozano2024ClinfoAI provides evidence-based healthcare QA by retrieving information from trusted sources like PubMed using the Entrez API, ensuring responses reflect current medical evidence. RALL Guo2024 employs Dense Passage Retriever to generate lay language explanations of medical concepts, addressing the growing need for accessible medical information. Self-BioRAG jeong2024improving implements the MedCPT retriever with self-reflection techniques to improve accuracy for complex clinical questions with multiple components, while WeiseEule Aftab2024 evaluates multiple retrievers (MedCPT, BioBERT, BioGPT) with Pinecone DB to handle diverse medical information needs across specialties. These medical QA systems collectively demonstrate RAG’s potential for healthcare information access, providing contextually relevant, evidence-based responses to complex medical queries for both clinical decision support and patient education.
@@
Diagnostic and treatment decision support systems assist clinicians in navigating complex clinical pathways through evidence retrieval and synthesis Sutton2020CDSS. They can help healthcare providers identify relevant evidence, recognize patterns, and identify appropriate interventions, ultimately improving diagnostic accuracy and treatment selection while reducing cognitive biases and diagnostic errors.
Several innovative implementations showcase RAG’s potential in clinical decision support across medical specialties. Expert-Guided LLMs Lammert2024ExpertLLM support specialized decision-making in precision oncology by retrieving relevant genomic information and matching it with treatment guidelines and clinical trial data. DR.KNOWS Gao2025 facilitates diagnostic prediction by connecting EHRs with medical knowledge repositories through knowledge graph-enhanced retrieval that models complex relationships between symptoms, diseases, and treatments. CliniqIR Abdullahi2024 combines MedCPT and BM25 retrieval to match patient symptoms with potential diagnoses, while SurgeryLLM ong2024surgeryllm delivers surgical decision support by retrieving relevant procedures, techniques, and complication data. RECTIFIER unlu2024retrieval employs OpenAI Text-Embedding with Faiss for clinical trial screening, helping match heart failure patients with appropriate clinical trials by analyzing eligibility criteria against patient data. By connecting patient-specific data with evidence-based medical knowledge, these RAG systems help clinical decision-making while reducing the cognitive burden on healthcare providers, potentially improving both efficiency and quality of care.
@@
Rare disease identification and management represents a particularly challenging healthcare domain due to limited literature, dispersed expertise, and complex symptom patterns MartinezDeMiguel2022RareDis. RAG systems are especially valuable in this context because they can aggregate and synthesize knowledge from diverse sources, helping clinicians recognize unusual patterns and connect them with appropriate diagnostic and management strategies for conditions they may encounter only rarely in clinical practice.
Recent research has produced several specialized rare disease applications leveraging different knowledge retrieval approaches. Zelin et al. Zelin2024RareDisease developed a RAG system specifically focused on rare disease diagnosis, connecting uncommon symptom patterns with rare disease knowledge drawn from specialized databases. RAG-HPO garcia2024deepphenotyping utilizes FastEmbed for automated deep phenotyping of rare genetic disorders, linking phenotypic observations with potential genetic causes through human phenotype ontology mapping. Yang et al. Yang2024RDguru created RDguru, an intelligent conversational agent that employs a multi-source fusion diagnostic model combining GPT-4, PheLR, and phenotype matching strategies to significantly improve diagnostic recall for rare diseases. A recent study kim2024phenotype has also evaluated the capabilities of LLMs and RAG for phenotype-driven gene prioritization in rare genetic disease diagnosis. These applications demonstrate how RAG systems can help bridge knowledge gaps in rare disease management by connecting dispersed information to support timely diagnosis and appropriate treatment strategies for conditions where clinical expertise is often limited.
@@
Automatic and semi-automatic report generation can help healthcare providers by significantly reducing the time they spend on documentation, allowing them to dedicate more time to patient care sloan2024automated. A critical area of report generation is radiology, which involves image analysis, document consultation, and data evaluation hamamci2024ct2rep. This process faces multiple challenges including ensuring factual accuracy, maintaining clinical relevance, minimizing hallucinations, and providing sufficient interpretability for clinical adoption alfarghaly2021automated.
RAG frameworks have been proposed to address these challenges by grounding language generation in relevant retrieved content, resulting in more accurate and trustworthy outputs xiong2024benchmarking. Recent works have demonstrated the effectiveness of RAG in this domain. For example, multimodally-aligned embeddings with general-domain LLMs such as GPT-4 ranjit2023retrieval, ensuring relevant content retrieval and mitigating hallucinations while allowing customization based on clinical intent. To enhance interpretability, some authors alam2024towards proposed a multi-agent RAG framework utilizing concept bottleneck models that guide report generation with clinically meaningful concept vectors, bridging the gap between performance and explainability. RAG pipelines bernardi2024report have also been used to improve image-text matching techniques, incorporating semantic segmentation and triple contrastive loss, leading to more accurate and fluent reports with reduced errors. Differently, LaB-RAG song2024lab employs label-boosted retrieval by converting medical images into descriptive text labels, which are then used for text-based retrieval and report generation, bypassing the need for direct image processing by the LLM. Lastly, some authors sun2024fact have proposed fact-aware multimodal retrieval strategies to address factual inconsistencies in generated reports by leveraging RadGraph-based factual mining and multimodal retrievers, ensuring high-quality reference retrieval and improved factual completeness.
@@
Genomic medicine involves analyzing genetic data to inform clinical decision-making. The rapid knowledge evolution and inherent complexity in variant interpretation within genomic medicine impose significant cognitive burdens on clinicians, particularly given the exponential expansion of genomic information and the challenges in ascertaining variant pathogenicity burton2012genomic.
Several innovative RAG applications address these challenges in genomic medicine through specialized knowledge retrieval. FAVOR-GPT li2024favorgpt utilizes OpenAI Text-Embedding with Weaviate DB to facilitate genome variant annotations, connecting genetic variants with clinical significance and supporting evidence from multiple databases. Lu et al. Lu2025BoostingGPT developed a genomics analysis system using Azure AI Search with keyword matching for more accurate genetic variant interpretation within clinical contexts. RAG-HPO garcia2024deepphenotyping connects clinical observations with genetic variants in rare disorders through sophisticated knowledge mapping, while JGCLLM Fukushima2025 implements GLuCoSE-base-ja vectors to provide genetic counseling support that translates complex genomic findings into clinically meaningful information. By connecting genomic findings with clinical knowledge through appropriate retrieval mechanisms, these RAG systems support the clinical use of genetic information for tasks such as disease inference Lu2025BoostingGPT as well as phenotypic and genotypic analysis garcia2024deepphenotyping, thereby enabling more personalized treatment decisions in the emerging era of genomic medicine.
@@
Personalized health management, a tailored healthcare approach, involves generating individualized health information and recommendations based on patient-specific characteristics, preferences, and health status Quazi2022HCI. This application mitigates a important healthcare challenge: the need for more effective and patient-centered care that optimizes treatment adherence, self-management, and health outcomes, in contrast to traditional standardized approaches that often fail to accommodate specific patient needs, literacy levels, and medical conditions.
Recent research has produced several personalized health management systems employing different knowledge retrieval strategies. HEALIE Kakalou2024 leverages knowledge graphs to generate personalized medical content incorporating both clinical knowledge and socio-economic factors relevant to individual patients. RISE wang2024diabetesrag implements OpenAI Text-Embedding with FAISS to provide personalized responses to diabetes-related inquiries, helping patients navigate complex self-management requirements. A Dual RAG System Lee2024 combines dense retrieval and BM25 with language-specific tokenizers to support diabetes management across multiple languages, expanding access to personalized health information for diverse populations. RAP azimi2025evaluation utilizes Amazon Titan Text Embeddings to provide personalized nutrition-related question answering, supporting dietary management for various health conditions. These personalized health management systems demonstrate RAG’s ability to bridge the gap between complex medical information and individual patient needs through contextually relevant knowledge retrieval, ultimately supporting better self-management and improved health outcomes through personalized guidance.
@@
Medical education and training systems support the development of healthcare professionals through evidence-based learning resources and simulated clinical experiences Lemmetty2006HIS. These systems help address workforce needs by enabling more efficient knowledge acquisition, supporting the development of clinical reasoning skills, and allowing for personalized learning experiences that adapt to individual learner needs and knowledge gaps.
ChatENT long2024chatent leverages Retrieval-Augmented Language Modeling (RALM) to support medical education in Otolaryngology-Head and Neck Surgery (OHNS). EyeGPT chen2024eyegpt transforms a general LLM into an ophthalmology-specialized assistant through prompt engineering, fine-tuning with eye-specific terminology, and RAG from ophthalmology textbooks, supporting both medical education and patient inquiries. Sevgi et al. sevgi2024medical developed targeted ophthalmology applications, including EyeTeacher for generating educational questions and EyeAssistant for clinical support. Both systems are enhanced by integrating clinical guidelines and peer-reviewed documents as retrieval sources. In nursing education, Zhao et al ZHAO2025104284 leveraged RAG and semantic search technology to answer questions from the Chinese Nursing Licensing Exam, showing the potential for specialized knowledge retrieval in professional certification. By connecting learning objectives with relevant medical knowledge through appropriate retrieval mechanisms, these RAG-based education systems help support healthcare practice and skill training.
@@
Clinical research application tools are designed to support clinical studies by enhancing efficiency, leveraging knowledge retrieval and synthesis to accelerate the development of medical knowledge and therapeutic innovations. ge2024leveraging.